A valószínűségszámítás II. gyakorlaton való részvétel

dátum:  II. 8. II. 15. II. 22. III. 1. III. 8. III. 22. IV. 5.

diákok:

Ferenczik Judit  +  +  +  +  +  +  +
Husiová Anikó  +  +  +  +  +  +  +
Kiss Andrea  +  +  +  +  +  −  +
Lörincz Tamás  −  +  +  +  +  −  +
Mikó Tamás  +  +  −  +  +  −  +
Nemoda Dóra  +  +  +  +  +  +  +
Pajkos Gabriella  +  −  +  +  +  +  +
Porvázsnyik Bettina Lívia  +  −  +  +  +  +  +
Szabó Lívia Dóra  +  +  −  +  +  +  +
Szűcs Tamás Gábor  +  +  +  +  +  −  −
Vadász Péter  −  −  +  −  −  −  −
Vadnai Zoltán  −  +  −  +  +  +  +

dátum:  IV. 12. IV. 19. V. 3. V. 10. V. 17.

diákok: Ferenczik Judit Husiová Anikó Kiss Andrea Lörincz Tamás Mikó Tamás Nemoda Dóra Pajkos Gabriella Porvázsnyik Bettina Lívia Szabó Lívia Dóra Szűcs Tamás Gábor Vadnai Zoltán